Re: pedAsPhys Fwd: Re: question?

Date view Thread view Subject view Author view Attachment view

From: Frank Geurts (geurts_at_rice.edu)
Date: Wed Jan 29 2003 - 15:38:52 EST


>> please remember to let me know (today or tomorrow)
>> if there are things i can do during the access wed-thurs...
>>
>
> noise measurements might be one of the first things that could be done
> .. it allows us to keep track of any changes but also should do it as a
> diagnostics tool (as in why we have this C-section problem).
>
> Here's something I don't completely understand: if i look at some
> localmon events I see that sometimes the last 8 TDC channels are 0. But,
> if the event is a strobe event the 0s are gone and I can see the
> common-stop signal again ... so the TDCs seem to be in place and get
> strobed. I am very sure from the last noise measurements i did that the
> B-section cables are hooked up (the values were reasonable enough to
> assume so and seem to have doubled compared to the measurements before).
> The last 8 TDC channels are from a set different discriminators all of
> which I set the threshold equal to that of the others- but that couldn't
> explain the overall shift in channels you mentioned earlier ...
>

just to make sure this triviality is not biting us: the numbering of
TOFr ADC and TDC channels is (by design) off by one ADC CAMAC module
(12channels). So for example in the tofr-range TDC#80 belongs to ADC#92.
 From your "before" plot it clearly shows this shift was missing, but
still doesn't explain the 8 channels (instead of 12!).

-frank


Date view Thread view Subject view Author view Attachment view

This archive was generated by hypermail 2.1.4 : Thu Jul 24 2003 - 00:39:34 EDT